A conducting single-turn circular loop with a tota
Answer #1

induced emf = d(phi_B)/dt

= b*2*t - c*3*t^2

= 2*b*t - 3*c*t^2


when, induced emf is maximum

d^2(phi_B)/dt^2 = 0

2*b - 3*c*2*t = 0

==> t = 2*b/6*c

= 2*13.5/(6*5.5)

= 0.818 s

so, maximum induced emf = 2*b*t - 3*c*t^2

= 2*13.5*0.818 - 3*5.5*0.818^2

= 11.05 volts <<<<<------------Answer